# Concept Complement Bottleneck Model for Interpretable Medical Image Diagnosis

### v1

@@ Abstract
Models based on human-understandable concepts have received extensive attention to improve model interpretability for trustworthy artificial intelligence in the field of medical image analysis. These methods can provide convincing explanations for model decisions but heavily rely on the detailed annotation of pre-defined concepts. Consequently, they may not be effective in cases where concepts or annotations are incomplete or low-quality. Although some methods automatically discover effective and new visual concepts rather than using pre-defined concepts or could find some human-understandable concepts via large Language models, they are prone to veering away from medical diagnostic evidence and are challenging to understand. In this paper, we propose a concept complement bottleneck model for interpretable medical image diagnosis with the aim of complementing the existing concept set and finding new concepts bridging the gap between explainable models. Specifically, we propose to use concept adapters for specific concepts to mine the concept differences and score concepts in their own attention channels to support almost fairly concept learning. Then, we devise a concept complement strategy to learn new concepts while jointly using known concepts to improve model performance. Comprehensive experiments on medical datasets demonstrate that our model outperforms the state-of-the-art competitors in concept detection and disease diagnosis tasks while providing diverse explanations to ensure model interpretability effectively.
{IEEEkeywords}
Concept Bottleneck Model, Medical Image Diagnosis, Interpretability, XAI.
@@ 1 Introduction
Interpretable deep learning models can be divided into post-hoc models and ante-hoc models [6]. Post-hoc models provide final explanations for model decisions after the model training process, such as gradient-based methods [7, 8] and counterfactual-based methods [9, 10]. Although post-hoc explanations are expected to provide strong evidence for model decisions in many applications [11, 12], these models are sensitive to input data, so their decisions are more easily influenced by changes in training data, indicating that the explanations are unreliable [13]. Ante-hoc models ensure inherent interpretability throughout the entire process of model training and inference. In the field of MIA, ante-hoc explainable models mainly include attention-based [14, 15], example-based [16, 17, 18] and concept-based models [19, 20, 21, 22].
@@
Among these models, concept-based models have caught the most eyes of researchers since they rely on human-understandable textual concepts or human-friendly visual concepts to explain model decisions [23]. Some studies focus on predicting concepts contained in the images and detailedly annotated by doctors or experts to make decision. For example, the Concept Bottleneck Model (CBM) [24] is one of the most representative works. CBM predicts concept scores by minimizing the concept classification cross-entropy loss. Then, it jointly or independently trains a classifier with a bottleneck layer to make decision in which step it allows externally adjusting the contribution of concepts to modify the final decision. Furthermore, many variants of CBM have been extensively studied in disease diagnosis based on X-ray, ultrasound and other medical images [19, 25]. Although CBM-based methods effectively enhance model interpretability, they require fully concept annotations of training data. However, in real-world scenarios, detailed annotations are extremely time-consuming and labor-consuming. There are some approaches that could automatically learn visual concepts [26, 27], but they finds concepts which are difficult to correspond to clinical concepts in a general way. With the development of Large Language Models (LLMs) or Visual Language Models (VLMs), although many researches could automatically discover and generate textual concepts for images by using LLMs or VLMs [28, 29], these concepts make more sense in the general field, but not so much in the medical field.
@@
Despite that the mentioned methods have achieved good interpretability and classification performance, they rely on annotated concepts completely or only use discovered concepts for decision-making. The former requires dense concept annotations and performs less effectively than black-box models, while the latter can easily deviate from real medical decision-making processes. Additionally, most of existing concept-based methods use the same image features for concepts without considering the differences among concepts. On the visual level, simpler concepts are easier to capture, but in general, simple concepts contribute less to model decisions, especially for samples that are difficult to diagnose. Therefore, it is unfair to directly use the same feature encoded by an image encoder to calculate concept scores.
@@
To address these issues, we propose an interpretable model named Concept Complement Bottleneck Model (CCBM) for medical image disease diagnosis, it aims to achieve fair concept learning, keep the explanations from existing concept annotations and discover additional concepts that could collaborate with pre-defined concepts to help the model approach the performance of black-box models. Specifically, our contributions are as follows: 1) Considering that one image encoder is hard to capture all concepts comprehensively, we configure a concept adapter for each concept to encode the most relevant part of the corresponding concept from the basic image feature. We also leverage Multi-Head Cross-Attention (MHCA) to calculate concept scores in their own channels to support fair concept learning and provide a flexible structure to help concept discovery. 2) We design a concept complement strategy to support unknown concept learning when jointly using known concepts, which could help the model approximate black-box performance by adding unnconcept adapters and the concept aggregator directly. 3) CCBM can provide diverse visual and textual explanations for model decision. Especially, benefit from independently concept feature extracting and scoring, we can visualize all known or unknown concepts more effectively. 4) Extensive experiments demonstrate the effectiveness of our method in the aspects of concept detection as well as disease diagnosis. Model interpretability is verified by quantitative and qualitative experiments in the aspects of faithfulness, label effectiveness and understandability.
@@ 2 Related Work
Model interpretability is a crucial issue, especially in the field of MIA. The decisions of deep learning models in this high-risk area directly relate to human life and health. Therefore, for doctors and patients, any model decision must be reasonably explained, which could promote humna trust model and make the next medical decision based on the model prediction. To enhance users’ confidence in computer-aided or automated diagnostic systems, concept-based interpretable models for disease diagnosis have been studied widely.
With the development of CBM, some research has been proposed to enhance the interpretability of disease diagnosis models based on the concept bottleneck. For instance, Yuksekgonul et al. [30] propose a Post-hoc Concept Bottleneck Model (PCBM) which flexibly obtains concept vectors from datasets with annotated concepts to generate concept subspace. Then, they map the feature of target domain data to the concept subspace to derive concept scores. Additionally, they also incorporate a residual network to enhance classification performance at the expense of model transparency. Chauhan et al. [19] introduce an interactive strategy that enables the model to seek input from human collaborators for labeling specific concepts, thereby enhancing the model’s accuracy. Furthermore, some studies leverage textual concept encoding to aid in concept identification in images. Patricio et al. [31] propose a two-stage skin disease diagnosis model. Initially, they deriv lesion segmentation masks by a trained semantic segmentation model to masking invalid features. Subsequently, they append a concept encoding layer after the feature extractor to pinpoint the presence of specific concepts in the image and generate corresponding scores for classifier training. During training, they minimize the consistency loss between final visual concept encoding and textual concept encoding. Bie et al. [32] present a skin disease diagnosis method based on multi-level alignment of image and textual concepts encoded by the pretrained LLM. They capture more detailed semantic information in the images through three levels of alignment: image-level, token-level, and concept-level. Similarly, they employ a multilayer perceptron with two fully connected layers (FCLs) to compute concept scores and diagnose skin diseases from aligned features.
@@ 3.1 Textual and Visual Concept Feature Extraction
We first extract the textual known concept embeddings and then specifically learn visual features through the concept adaptersvin this step.
@@ 3.1.2 Encoding visual known concepts independently
We set up concept adapters to extract key features from a shared image encoding for each concept, considering their unique differences. Detailedly, we select a Convolutional Neural Network (CNN) $E_{I}$ as image encoder to map the input image data $\mathcal{X}$ including $N$ samples from $n_{c}$ classes to the fundamental feature space: $E_{I}:$ $\mathcal{X}\rightarrow\mathbb{R}^{d}$, where $d$ represents the feature dimension. After that, we set $n_{k}$ concept adapters $C_{i}$ $(i=1,2,...,n_{k})$ to extract specific concept features from the fundamental features to be the quaries $Q$ of the following cross-attention module. Intuitively, we hope the fundamental features extracted by the image encoder can include all of crucial features for all different concepts, so we had better to choose a strong image encoder. For simplicity, in our setting, each concept adapter is set as a FCL that maps the $d$-dimension fundamental feature space to the ${d_{k}}$-dimension concept subspace. It is worth noting that we need to ensure that the subspace dimensions of all concept adapters are consistent with the subspace dimension of text encoder. We fomulate the image encoder and concept adapters as follows:
@@ 3.2 Concept Score Prediction
Through the encoders and concept adapters, we can obtain the image features of all known concepts and the embedings of all textual known concepts. At this step, we use the multi-head cross-attention module to obtain the attention score. For all known concepts, we use the embeddings of the textual concept as the keys and values, and the outputs of the concept adapters as the queries to calculate the attention weights where each output is a query. All textual concept embeddings are represented as the matrix $K\in\mathbb{R}^{m\times d_{k}}$ and $V\in\mathbb{R}^{m\times d_{k}}$, and all visual concept embeddings are denoted as the matrix $Q\in\mathbb{R}^{m\times d_{c}}$. In this paper, we only fomulate the expression of single-head cross-attention output:
$A(Q,K)=softmax\left(\frac{QK^{T}}{\sqrt{d_{c}}}\right),$ (2)
@@
where $A\in\mathbb{R}^{m\times d_{k}}$ is the attention map matrix whose elements represents the weights of different concept pairs and $A_{w}\in\mathbb{R}^{m\times d_{k}}$ is the weighted sum of all concepts. According to the setting of the concept adapters, we can not average these attention as final features to calculate concept scores by a FCL but need to aggregate them in another way to get concept scores independently. Specifically, we can get the attention $A_{w}(Q_{i},K,V)$ for the $i$-th concept. Furthermore, we need to calculate the concept scores based on these attention weights. Different from the previous bottleneck models who directly use a FCL to project the common feature to get the concept scores, we could calculate these concept scores independently by any score calculation module based on their specific concept features. In our model, the aggregator for known concepts includes $n_{k}$ FCLs, and we use each FCL $f_{i}$ to project the attention weights to get the concept score for each concept. Hence, the concept scores are calculated as follows:
$S_{i}=f_{i}(A_{w}(Q_{i},K,V))\in\mathbb{R}^{n_{k}},i=1,2,...,n_{k},$ (4)
@@ 3.3.1 Explainable Diagnosis Decision using Pre-defined Concept Set
where $S=[S_{1},S_{2},...,S_{n_{k}}]$ is the concept score vectors of input images and $\hat{Y}$ is the final disease prediction.
@@
and as for the latter, if the concept annotation is scored but not binary, we use the Mean Square Error (MSE) loss $L_{mse}^{cep}$:
$\mathcal{L}_{cep}=\sum_{i=1}^{N}\sum_{j=1}^{n_{k}}(c_{ij}-s_{ij})^{2},$ (9)
@@
Table 1: Dataset Details in Our Experiments.
Dataset Type Used Concepts Categories
Derm7pt Dermoscopic Atypical Pigment Network (APN), Blue Whitish Veil (BWV), Irregular Vascular Structures (IVS), Irregular Pigmentation”(IP), Irregular Streaks (IS), Irregular Dots and Globules (IDaD), Regression Structures (RS) Nevus Melanoma
Skincon Dermoscopic PAPule (PAP), PLAque (PLA), PUStule (PUS), BULla (BUL), PATch (PAT), NODule (NOD), ULCer (ULC), CRUst (CRU), EROsion (ERO), ATRophy (ATR), EXUdate” (EXU), TELangiectasia (TEL), SCALe (SCAL), SCAR” (SCAR), FRIable (FRI),Dome-SHaped (DSH), Brown-Hyperpigmentation (BrH), White-Hypopigmentation (WhH), PURple (PUR), YELlow” (YEL), BLAck” (BLA) and ERYthema (ERY) Malignant Benign Non-neoplastic
BrEaST Ultrasound Irregular SHape (IRS), Not Circumscribed Margin (NCM), Hyperechoic or Heterogeneous Echogenicity (HoHE), Posterior Features (PF), Hyperechoic Halo (HH), CALcifications (CAL), Skin Thickening (ST) Malignant Benign
LIDC-IDRI CT SUBtlety (SUB), InternalSTructure (IST), CALcification (CAL), SPHericity (SPH), MARgin (MAR), LOBulation (LOB), SPIculation (SPI), TEXture (TEX) Nevus Melanoma
@@ 3.3.2 Concept Complement Strategy
Although existing models based on concept bottleneck can maintain high classification performance while providing interpretability by training on concept detection tasks and classification tasks, the model performance needs to be balanced between concept detection and classification tasks. In addition, there is still a gap between explainable model and black-box models in classification tasks. In order to reduce this performance gap while maintaining model transparency, we propose a concept complement strategy to learning unknown concepts which are helpful to diagnosis.
The concept complement strategy includes two parts: concept complement and concept score adjustment. Concept complement means that we add $n_{u}$ additional unknown concept adapters which are set as $n_{u}$ FCLs projecting features from $d$-dimension to $d_{u}$-dimension space to learn $n_{u}$ new concepts. The general processing of unknown concepts is similar to that of known concepts in the other steps. However, we have no textual representions of these new concepts. We set $n_{u}$ learnable embedding vectors in the model to uniquely represent the learned unknown concepts which will be used as the keys $K_{j}^{u}$ and values $V_{j}^{u}$ of unknown concept cross-attention module. If we denote the visual encoding of unknown concepts from concept adapters as $Q_{j}^{u}\in\mathbb{R}^{n\times d_{u}}$, the concept scores of unknown concepts are calculated as:
@@
Table 2: Quantitative results on disease diagnosis and concept detection tasks with the state-of-the-art comparison methods and black-box models. The results are shown as the mean $\pm$ std of five-fold cross-validation experiment.

### v2

@@ Abstract
Models based on human-understandable concepts have received extensive attention to improve model interpretability for trustworthy artificial intelligence in the field of medical image analysis. These methods can provide convincing explanations for model decisions but heavily rely on the detailed annotation of pre-defined concepts. Consequently, they may not be effective in cases where concepts or annotations are incomplete or low quality. Although some methods automatically discover effective and new visual concepts rather than using pre-defined concepts or could find some human-understandable concepts via large language models, the discovered concepts are prone to veering away from medical diagnostic evidence or are challenging to understand. In this paper, we propose a concept complement bottleneck model for interpretable medical image diagnosis with the aim of complementing the existing concept set and finding new concepts to bridge the gap between explainable models and black-box models. Specifically, we propose to use concept adapters for specific concepts to mine the concept differences and score concepts in their own attention channels to support fairer concept learning. Then, we devise a concept complement strategy to learn new concepts while jointly using known concepts to improve model performance. Comprehensive experiments on medical datasets demonstrate that our model outperforms the state-of-the-art competitors in concept detection and disease diagnosis tasks while providing diverse explanations to ensure model interpretability effectively.
keywords:
Concept Bottleneck Model , Medical Image Diagnosis , Interpretability , XAI
journal: Medical Image Analysis
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
The proposed Concept Complement Bottleneck Model can leverage concept adapters and multi-head cross-attention to improve model performance on disease diagnosis and concept detection tasks.
Adding a concept adapter for each concept supports fairer concept learning so that the performance of concept detection is more balanced across different concepts.
Concept complement strategy can learn unknown concepts by jointly training with known concepts to help bridge the performance gap between explainable models and black-box models.
Extensive experiments demonstrate the effectiveness of concept complement bottleneck model in concept detection and disease diagnosis tasks while providing diverse visual and textual explanations for model decision.
@@ 1 Introduction
Interpretable deep learning models can be divided into post-hoc models and ante-hoc models [6]. Post-hoc models provide final explanations for model decisions after the model training process, such as gradient-based methods [7, 8] and counterfactual-based methods [9, 10]. Although post-hoc explanations are expected to provide strong evidence for model decisions in many applications [11, 12], their model decisions are sensitive to input data, indicating that the explanations are unreliable [13]. Ante-hoc models ensure inherent interpretability throughout the entire process of model training and inference [14]. In the field of MIA, ante-hoc explainable models mainly include attention-based [15, 16], example-based [17, 18, 19] and concept-based models [20, 21, 22, 23].
@@
Concept-based ante-hoc explainable models have attracted widespread attention due to their ability to explain model decisions using human-friendly textual or visual concepts [24]. Some studies focus on predicting concepts contained in the images and detailedly annotated by doctors or experts to make decision. For example, the Concept Bottleneck Model (CBM) [25] is one of the most representative works. CBM predicts concept scores by minimizing the concept classification cross-entropy loss. Then, it jointly or independently trains a classifier with a bottleneck layer to make decision in which step it allows externally adjusting the contribution of concepts to modify the final decision. Furthermore, many variants of CBM have been extensively studied in disease diagnosis based on X-ray, ultrasound and other medical images [20, 26]. Although CBM-based methods effectively enhance model interpretability, they require fine-grained concept annotations of training data. However, in real-world scenarios, obtaining detailed annotations are extremely time-consuming and labor-consuming. There are some approaches that could automatically learn visual concepts [27, 28], but the discovered concepts are difficult to correspond to clinical concepts in a general way. With the development of Large Language Models (LLMs) and Visual Language Models (VLMs), although many researches could automatically discover and generate textual concepts for images by using LLMs or VLMs [29, 30], these concepts make more sense in the general field, but not so much in the medical field.
@@
Despite that the mentioned methods have achieved good interpretability, they rely on annotated concepts completely or only use discovered concepts for decision-making. The former requires dense concept annotations and there is a performance gap between them and the black-box models [14], while the latter can easily deviate from real medical decision-making processes. Additionally, most of existing concept-based methods use the same image features for all concepts without considering the differences among concepts. On the visual level, simpler concepts are easier to capture, but in general, simple concepts contribute less to model decisions, especially for samples that are difficult to diagnose. Therefore, it is unfair to directly use the same feature encoded by an image encoder to calculate concept scores.
@@
To address these issues, we propose an interpretable model named Concept Complement Bottleneck Model (CCBM) for medical image disease diagnosis. It aims to achieve fair concept learning, keep the explanations from existing concept annotations and discover additional concepts that could collaborate with pre-defined concepts to help the model approach the performance of black-box models. Specifically, our contributions are as follows: 1) Considering that one image encoder is hard to capture all concepts comprehensively, we configure a concept adapter for each concept to encode the most relevant part of the corresponding concept from the basic image feature to support fair concept learning and capture the regions most related to the concepts via Multi-Head Cross-Attention (MHCA). 2) We design a concept complement strategy to support unknown concept learning when jointly using known concepts, which could help the model approximate black-box performance by adding concept adapters and the concept aggregator directly. 3) CCBM can provide diverse visual and textual explanations for model decision. Especially, benefit from independent concept feature extracting and scoring, we can visualize all known or unknown concepts more effectively. 4) Extensive experiments demonstrate the effectiveness of our method in the aspects of concept detection as well as disease diagnosis. Model interpretability is verified by quantitative and qualitative experiments in the aspects of faithfulness, label effectiveness and understandability.
@@ 2 Related Work
Model interpretability is a crucial issue, especially in the field of MIA. The decisions of deep learning models in this high-risk area directly relate to human life and health. Therefore, for doctors and patients, any model decision must be reasonably explained, which could promote human trust model and make the next medical decision based on the model prediction. To enhance users’ confidence in computer-aided or automated diagnostic systems, concept-based interpretable models for disease diagnosis have been studied widely.
With the development of CBM, some research has been proposed to enhance the interpretability of disease diagnosis models based on the concept bottleneck. For instance, Yuksekgonul et al. [31] propose a Post-hoc Concept Bottleneck Model (PCBM) which flexibly obtains concept vectors from datasets with annotated concepts to generate concept subspace. Then, they map the feature of target domain data to the concept subspace to derive concept scores. Additionally, they also incorporate a residual network to enhance classification performance at the expense of model transparency. Chauhan et al. [20] introduce an interactive strategy that enables the model to seek input from human collaborators for labeling specific concepts, thereby enhancing the model’s accuracy. Furthermore, some studies leverage textual concept encoding to aid in concept identification in images. Patricio et al. [32] propose a two-stage skin disease diagnosis model. Initially, they derive lesion segmentation masks by a trained semantic segmentation model to masking invalid features. Subsequently, they append a concept encoding layer after the feature extractor to pinpoint the presence of specific concepts in the image and generate corresponding scores for classifier training. During training, they minimize the consistency loss between final visual concept encoding and textual concept encoding. Bie et al. [33] present a skin disease diagnosis method based on multi-level alignment of image and textual concepts encoded by the pretrained LLM. They capture more detailed semantic information in the images through three levels of alignment: image-level, token-level, and concept-level. Similarly, they employ a multilayer perceptron with two fully connected layers (FCLs) to compute concept scores and diagnose skin diseases from aligned features.
@@ 3.1 Textual and Visual Concept Feature Extraction
We first extract the textual known concept embeddings and then specifically learn visual features through the concept adapters in this step.
@@ 3.1.2 Encoding Visual Known Concepts Independently
We set up concept adapters to extract key features from a shared image encoding for each concept, considering their unique differences. Detailedly, we select a Convolutional Neural Network (CNN) $E_{I}$ as the image encoder to map the input image data $\mathcal{X}$ including $n$ samples from $n_{c}$ classes to the fundamental feature space: $E_{I}:$ $\mathcal{X}\rightarrow\mathbb{R}^{d}$, where $d$ represents the feature dimension. After that, we set $n_{k}$ concept adapters $C_{i}$ $(i=1,2,...,n_{k}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{k}}$ to extract specific concept features from the fundamental features to be the queries $Q$ of the following cross-attention module. Intuitively, we hope the fundamental features extracted by the image encoder can include all of crucial features for all different concepts, so we had better to choose a strong image encoder. For simplicity, in our setting, each concept adapter is set as a FCL that maps the $d$-dimension fundamental feature space to the ${d_{k}}$-dimension concept subspace. It is worth noting that we need to ensure that the subspace dimensions of all concept adapters are consistent with the subspace dimension of text encoder. We fomulate the image encoder and concept adapters as follows:
@@ 3.2 Concept Score Prediction
Through the encoders and concept adapters, we can obtain the image features of all known concepts and the embedings of all textual known concepts. In this step, we use the multi-head cross-attention module to obtain the attention score. For all known concepts, we use the embeddings of the textual concept as the keys and values, and the outputs of the concept adapters as the queries to calculate the attention weights where each output is a query. All textual concept embeddings are represented as the matrix $K\in\mathbb{R}^{n_{k}\times d_{k}}$ and $V\in\mathbb{R}^{n_{k}\times d_{k}}$, and all visual concept embeddings are denoted as the matrix $Q\in\mathbb{R}^{n_{k}\times d_{k}}$. In this paper, we only fomulate the expression of single-head cross-attention output:
$A(Q,K)=softmax\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right),$ (2)
@@
where $A\in\mathbb{R}^{n_{k}\times n_{k}}$ is the attention map matrix whose elements represent the weights of different concept pairs and $A_{w}\in\mathbb{R}^{n_{k}\times d_{k}}$ is the weighted sum of all concepts. According to the setting of the concept adapters, we can not average these attention as final features to calculate concept scores by a FCL but need to aggregate them in another way to get concept scores independently. Specifically, we can get the attention $A_{w}(Q_{i},K,V)$ for the $i$-th concept. Furthermore, we need to calculate the concept scores based on these attention weights. Different from the previous bottleneck models which directly use a FCL to project the common feature to get the concept scores, we could calculate these concept scores independently by any score calculation module based on their specific concept features. In our model, the aggregator for known concepts includes $n_{k}$ FCLs, and we use each FCL $f_{i}$ to project the attention weights to get the concept score for each concept. Hence, the concept scores are calculated as follows:
$S_{i}=f_{i}(A_{w}(Q_{i},K,V)),i=1,2,...,n_{k},$ (4)
@@ 3.3.1 Explainable Diagnosis Decision using Pre-defined Concept Set
where $S=[S_{1},S_{2},...,S_{n_{k}}]$ is the concept score vectors of input images, $n_{c}$ is the number of classes and $\hat{Y}$ is the final disease prediction.
@@
and as for the latter, if the concept annotation is scored but not binary, we use the Mean Square Error (MSE) loss $\mathcal{L}_{mse}^{cep}$:
$\mathcal{L}_{cep}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}(c_{ij}-s_{ij})^{2},$ (9)
@@ 3.3.2 Concept Complement Strategy
Although existing models based on concept bottleneck can maintain high classification performance while providing interpretability by training on concept detection tasks and classification tasks, the model performance needs to be balanced between concept detection and classification tasks. In addition, there is still a gap between explainable models and black-box models in classification tasks. In order to reduce this performance gap while maintaining model transparency, we propose a concept complement strategy to learn unknown concepts which are helpful to diagnosis.
The concept complement strategy includes two parts: concept complement and concept score adjustment. Concept complement means that we add $n_{u}$ additional unknown concept adapters which are set as $n_{u}$ FCLs projecting features from $d$-dimension to $d_{u}$-dimension space to learn $n_{u}$ new concepts. The general processing of unknown concepts is similar to that of known concepts in the other steps. However, we have no textual representions of these new concepts. We set $n_{u}$ learnable embedding vectors in the model to uniquely represent the learned unknown concepts which will be used as the keys $K_{j}^{u}$ and values $V_{j}^{u}$ of unknown concept cross-attention module. If we denote the visual encoding of unknown concepts from concept adapters as $Q_{j}^{u}\in\mathbb{R}^{n_{u}\times d_{u}}$, the concept scores of unknown concepts are calculated as:
@@
4 Experiments
Table 1: Concepts in Our Experiments.
Dataset Used Concepts
Derm7pt Atypical Pigment Network (APN), Blue Whitish Veil (BWV), Irregular Vascular Structures (IVS), Irregular Pigmentation(IP), Irregular Streaks (IS), Irregular Dots and Globules (IDaD), Regression Structures (RS)
Skincon PAPule (PAP), PLAque (PLA),PUStule (PUS), BULla (BUL), PATch (PAT), NODule (NOD), ULCer (ULC), CRUst (CRU), EROsion (ERO), ATRophy (ATR), EXUdate (EXU), TELangiectasia (TEL), SCALe (SCAL), SCAR (SCAR), FRIable (FRI),Dome-SHaped (DSH), Brown-Hyperpigmentation (BrH), White-Hypopigmentation (WhH), PURple (PUR), YELlow (YEL), BLAck (BLA), ERYthema (ERY)
BrEaST Irregular SHape (IRS), Not Circumscribed Margin (NCM), Hyperechoic or Heterogeneous Echogenicity (HoHE), Posterior Features (PF), Hyperechoic Halo (HH), CALcifications (CAL), Skin Thickening (ST)
LIDC-IDRI SUBtlety (SUB), InternalSTructure (IST), CALcification (CAL), SPHericity (SPH), MARgin (MAR), LOBulation (LOB), SPIculation (SPI), TEXture (TEX)
4.1 Implementation Details
As for dermoscopic image datasets Derm7pt and Skincon, we follow PCBM [31] and choose the trained Inception-v3 model [39] as the image encoder, and we use pretrained ResNet50 as the image encoder for other two datasets. The frozen pre-trained ClinicalBERT [40] is utilized as the text encoder. In our experiments, the concept adapters and aggregators are set as FCLs. Additionally, we train the model for 300 epochs at most using the Adam optimizer and we stop the training process if there is no significant change in training loss. We use the grid method to choose model hyperparameters. For the Derm7pt and BrEaST dataset, the hyperparameters are set as $\lambda_{1}=0.2,\lambda_{2}=10$. For the Skincon dataset, $\lambda_{1}=0.1,\lambda_{2}=5$, and $\lambda_{1}=0.5,\lambda_{2}=10$ for LIDC-IDRI dataset. As for the number of unknown concepts, we can set $n_{u}$ as the number of categories of datasets, since intuitively we need at least the same number of data features as the sample categories to distinguish all categories. To evaluate model performance, we use the Area Under Curve (AUC), ACCuracy (ACC) and F1-score as disease diagnosis evaluation metrics, and the first two metrics are also used to evaluate the concept detection tasks of Derm7pt, Skincon and BrEaST. Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) are used for evaluating the concept Regression task of LDIC-IDRI dataset. All of mean and standard deviation results are obtained by five-fold cross-validation in our experiments.
4.2 Comparison Algorithms
4.3 Experimental Results and Analysis
4.3.1 Disease Diagnosis and Concept Detection
Table 2: Quantitative results on disease diagnosis and concept detection tasks with comparison methods and black-box models. The results are shown as the mean $\pm$ std of five-fold cross-validation experiment.